# On misconceptions about the Brier score in binary prediction models

### v3

@@ 2.1. Key properties about the BS
    Optimal predictions and true probabilities: The (unique) optimal predictions minimising the expected Brier score are the true outcome probabilities, i.e., $p_{i}=q_{i}$, as the Brier score is a strictly proper scoring rule [9, Theorem 1].
@@
    Expectation of Brier score for $n=1$: Knowing the true probability, we can calculate the expectation. For $n=1$, let $Y_{1}\sim Bern(q_{1}),q_{1}\in[0,1]$ with prediction $p_{1}\in[0,1]$. Then the expectation of Brier score is given as
    (2.1) $g(p_{1},q_{1}):=\mathbb{E}_{Y_{1}\sim Bern(q_{1})}[BS(p_{1},Y_{1})]=p_{1}^{2}-2p_{1}q_{1}+q_{1}.$
    For the interested reader, the calculation is given in Appendix B.1. The optimal prediction minimising (2.1) is $p_{1}=q_{1}$. We next go through two key cases
@@
      • if $q_{1}=1/2=p_{1}$, the expected Brier score is $f(1/2)=1/4$ corresponding to the maximum of (2.2),
@@
      The expectation of the Brier score as a function of $q_{1}$, when $p_{1}=q_{1}$, i.e., (2.2), looks as follows:
@@
      • the expectation of the Brier score under perfect prediction but increased true probability, i.e., $\tilde{p}_{1}:=p_{1}+\varepsilon=\tilde{q}_{1}=q_{1}+\varepsilon$, and taking the difference
        (2.3) $g(q_{1}+\varepsilon,q_{1}+\varepsilon)-g(q_{1},q_{1})=\varepsilon(1-2q_{1}-\varepsilon)$
        Note that the difference depends strongly on $q_{1}$.
@@
        (2.4) $g(q_{1}+\varepsilon,q_{1})-g(q_{1},q_{1})=\varepsilon^{2}$
        Note that the difference depends does not depend on $q_{1}$, and the same holds for $\tilde{p}_{1}:=p_{1}-\varepsilon$. The calculations are in Appendix §$\S$ B.2.
    • From the previous point, we conclude that if we slightly predict wrong, the costs are almost inexistent, e.g. if the true probability is $q_{1}$ but we predicted $p_{1}=q_{1}+0.1$, the expected difference between the Brier score of the perfect prediction and the slightly wrong one is $0.01$ by (2.4). However, if the true probability is changed e.g. from $q_{1}=0.1$ to $\tilde{q}_{1}=0.2$ and we have perfect prediction, the expected difference of the BSs is $0.07$ by (2.3).
    • Furthermore, we note the following nice order preservation property from (2.4): If $p_{1},\tilde{p_{1}}\in[0,1]$ such that $|p_{1}-q_{1}|<|\tilde{p_{1}}-q_{1}|$, then
@@
      This property is known as effective in the literature on scoring rules, cf., e.g., [17].
@@
    Dependence of expectation and distribution of Brier score on true probabilities: The Brier score with several observations as in (1.1) consists of the mean value of the one-dimensional BS’s. The expectation of Brier score in (1.1) and its distribution depend on the true outcome probabilities $q_{i}$, which are generally unknown, even when these are used as predictions. Based on (2.2) or Figure 1 its follows e.g. that when the $q_{i}$ are mostly concentrated around zero and one, then the expectation of the Brier score under perfect predictions (1.1) is close to zero. On the other hand, when they are mostly concentrated around e.g. $0.5$, the expectation of the Brier score (1.1) with perfect predictions is close to $0.25$.
@@
    Unobservability of true probabilities in clinical data: In practice, the true probability of an event occurring for an individual patient is not directly observable. Each patient is unique, and we only observe whether the event occurs or not (i.e., a binary outcome). Note however, that in simulation settings one can compare the true probabilities to estimated probabilities of regression or ML models, i.e. as done here [39, 13].
In particular we note that although individual-level true probabilities will remain unknown, we can approximate their average value across a population by calculating the mean of observed outcomes in overall, or in similar patient groups. This provides a practical way to assess model calibration in clinical settings via calibration in the large (CIL) [7, 38]
(2.5) $CIL(p,y)=\frac{1}{n}\sum_{i=1}^{n}p_{i}-\frac{1}{n}\sum_{i=1}^{n}y_{i}.$
2.2. Notable mathematical features about the BS
Below, we summarise basic mathematical properties of the Brier score for the understanding of the reader. Mathematical details and arguments are given in Appendix B.3.
@@ 2.2. Notable mathematical features about the BS
    Typical Brier score with perfect prediction for large $n$: Assume the true probabilities $q_{i}$ itself are realisations of random variables $Q_{i}\sim F$. Then, by the law of large numbers(LLN) the Brier score (1.1) for $n$ big roughly equals $\mathbb{E}[BS(Q_{1},Y_{1})]$, and probabilities for deviation from this value can be calculated via the central limit theorem (CLT). Hence $\mathbb{E}[BS(Q_{1},Y_{1})]$ roughly equals $BS((q_{1},\cdots,q_{n}),(y_{1},\cdots,y_{n}))]$ for $n$ large, and $\bar{y}$ roughly equals $\bar{q}$. More detail on this perspective is in Appendix §$\S$ A.
2.3. Main takeaways of the observations on the BS
To summarise our previous points, we conclude that an observed Brier score is a function of
  1. (I)
    the underlying true probabilities (the $q_{i}$s),
  2. (II)
  3. (III)
The influence of the randomness decreases with $n$ (in idealised proper settings by LLN and CLT), and the expectation of the Brier score can be seen as the long term average, which represent typical values if $n$ is big.
2.4. Means of understanding the BS
  • Expectation of the BS: We will analyze the expected value the Brier score (1.1) takes, which will help to illustrate typical observed values of the BS.
  • Simulation-Based Evaluations: We assess the BS’s behavior by simulating Bernoulli outcomes with different probability distributions for $q_{i}$ and different sample sizes. This approach intuitively explores properties in both realistic and idealized medical prediction scenarios.We will use the following elements in our simulations, where we give more detail in the Appendix §$\S$ E via the Aims, Data-generating mechanism, Estimands and other targets, Methods, Performance measures (ADEMP) framework [29].
    • Sample size $n$:. Settings considered: $n\in\{300,1000\}$.
    • True distribution: The $q_{i}\in[0,1]$ used to simulate $Y_{i}\sim Bern(q_{i})$. The $q_{i}$ itself are considered as realisations of random variables.
    • Predictor distribution: The $p_{i}\in[0,1]$ used in (1.1). These are considered as functions of $q_{i}$ with potentially random error.
    • Estimand: Includes median Brier score and $2.5\%$ resp. $97.5\%$ quantiles.
3. Misconceptions
  • A Brier score of 0 means a perfect model. A Brier score of 0 implies perfect alignment between predicted probabilities and observed outcomes. While the true probabilities are unknown, they lie within $[0,1]$, and we usually expect that a majority is in $(0,1)$. Hence, rather than signalling a perfect model, an observed Brier score of 0 potentially indicates a serious error. Given the binary outcomes, this would imply predictions were exclusively 0 or 1, contradicting the goal of probability estimation.
  • A perfect model has Brier score 0. Note that under the assumption that at least one of the true probabilities is in $(0,1)$, a Brier score of 0 is mathematically impossible perfect predictions (see Appendix C). Consider expected values of a perfect model. If $n=1$, and say $q_{1}=0.1$, this expectation is not equal to zero for the perfect model by (2.2), nor can an observed Brier score be equal to zero. If we evaluate several perfect predictions, the corresponding expected value of the Brier score is the mean over (2.2) which is not zero (unless for all $i$, $q_{i}\in\{0,1\}$).
To illustrate this and for later reference, we present simulations with perfect predictions, showing that even ideal models do not yield a Brier score of 0. In fact, expected Brier scores for a perfect model can be notably high.
True Predictor n Median of BS
distribution distribution (2,5%-&97,5%- Q)
Unif(0,1) Perfect 300 $0.167(0.144,0.19)$
Unif(0,1) Perfect 1000 $0.167(0.154,0.179)$
Unif(0,1) $+Unif(-0.2,0.2)$ 300 $0.179(0.156,0.203)$
Unif(0,1) $+Unif(-0.2,0.2)$ 1000 $0.179(0.165,0.193)$
Smoke prediction Perfect 300 $0.168(0.145,0.191)$
Smoke prediction Perfect 1000 $0.168(0.156,0.181)$
Smoke prediction Biased +0.1 300 $0.178(0.161,0.196)$
Smoke prediction Biased +0.1 1000 $0.178(0.169,0.188)$
Osteo prediction Perfect 300 $0.054(0.037,0.074)$
Osteo prediction Perfect 1000 $0.054(0.045,0.065)$
Osteo prediction Biased +0.1 300 $0.064(0.05,0.08)$
Osteo prediction Biased +0.1 1000 $0.064(0.057,0.073)$
Beta(2,2) Perfect 300 $0.2(0.18,0.22)$
Beta(2,2) Perfect 1000 $0.2(0.189,0.211)$
Table 2. Comparison median of Brier scores under different simulation settings
Prediction models can be compared across datasets through Brier score, but caution is essential. While a lower Brier score for one prediction model compared to another may suggest better predictive performance, this can be deceptive. Below, we present three scenarios with decreasing levels of caution, each illustrating why direct comparison of BSs may be misleading.
3.2.1. Comparing BSs Across Different Datasets with different incidences:
If the true outcome distributions differ, even perfect models will yield different Brier score distributions and expectations. Thus, cross-dataset comparisons may lack meaningful insights. Observing different incidences indicates that the true outcome distributions differ, making the Brier score comparison unreliable. The Brier score is a function of the true underlying probabilities, the predictions and some randomness, with the true probabilities having a big influence. Hence a lower Brier score in prediction model $A$ when compared to prediction model $B$ does not necessarily indicate a superior model in $A$ with more accurate predictions, as the expectation of Brier score depends on the true probabilities (BV). Compare, e.g. the osteoporosis prediction with $0.1$ bias to the perfect smoke prediction model from Table 5. The biased osteoporosis prediction has median Brier score $0.064$, while perfect smoke prediction has median Brier score $0.168$, nonetheless the perfect model is obviously better.
3.2.2. Comparing BSs Across Different Datasets with very similar or same incidences (also known as class imbalance) is meaningful.
Some recommend to compare Brier score of two prediction models on datasets with similar class imbalances [11]. However, the true probability distribution is unobservable, and by (BV) the true values $q_{i}$ determine the expectation of the Brier score. Thus, even with identical class imbalances, we cannot assume that the true probability distributions are comparable and the same caution as in the previous point should be exercised in interpreting such comparison. Descriptions in population characteristics can indicate similarities and differences of true probabilities. Compare the example with underlying distribution $Unif(0,1)$ where the prediction has error $+Unif(-0.2,0.2)$ to the the example with $Beta(5,5)$ under perfect predictions from Table 5. Both $Unif(0,1)$ and $Beta(5,5)$ give expected incidences of $50\%$. The Brier score of predictions with errors but underlying distribution $Unif(0,1)$ has median Brier score $0.179$, while perfect prediction for underlying $Beta(5,5)$ has median Brier score $0.200$, nonetheless the perfect model is obviously better.
3.2.3. Comparing Models on the same dataset via Brier score:
Even when evaluating two models on the same dataset, a lower Brier score in one does not always strictly imply that this model is better. There are two potential issues:
  1. (I)
    It is possible that the model that is worse has a better score by pure chance. However, with higher sample sizes this becomes more unlikely by the CLT (MIII).
  2. (II)
    The way in which the Brier score in expectation judges which model is better might not align with how a human might judge. In expectation, a change of the prediction from the perfect prediction $q_{i}$ to $q_{i}+\varepsilon$ or $q_{i}-\varepsilon$ gives the same result by (2.4). However, for humans the direction can matter as well, particularly for $q_{i}$ close to zero or one, as observed in [27]. Consider the example with one observation from [27, §$\S$ 3] where the probability of an event is $P(Y_{1}=1)=\frac{1}{10}$ (i.e., 10% chance), and compare two models:
    • Model 1: $p_{1}=0$ (predicting the event will never occur)
    Using (2.1), model 1 has an expected Brier score of $0.1$, as compared to $0.1125$ for model 2. As model 1 predicts a zero probability for the observation that has actually a 10% probability, at least intuitively, 25% might seem better for human judgement.
3.3. Misconception #3: A low Brier score indicates good calibration.
A low Brier score does not necessarily indicate good calibration of a model. As we have repeated, the Brier score is a function of the true underlying probabilities, the predictions and some randomness, with the true probabilities having a big influence. Hence it is possible that we have perfect predictions, but low or high Brier score due to the underlying probabilities, or similary not so accurate predictions but a Brier score that is low or high Brier score due to the underlying probabilities, see (BIV). As in expectation a change of the prediction from the perfect prediction $q_{i}$ to $q_{i}+\varepsilon$ or $q_{i}-\varepsilon$ gives the same result by (2.4), miscalibration where errors tend to go mostly in one direction are not punished more. Hence in the examples below from Table 3 we observe that median Brier score from the prediction biased by $0.1$ are the same as the one from the prediction with $+(1-2Bern(1/2))\cdot 0.1$ which has prediction errors that are as often and as much higher as they are lower than the true probability..
Assessing calibration should be done using additional metrics like CIL, calibration curves or similar evaluation components [7, 38]. Therefore, while the Brier score is a useful measure, it should not be the sole criterion for evaluating model performance.
True Predictor n Median CIL Median Brier score
distribution distribution (2,5%-&97,5%- Q) (2,5%-&97,5%- Q)
Unif(0,1) Perfect 1000 $0(-0.025,0.025)$ $0.166(0.154,0.179)$
Unif(0,1) $+0.1$ 1000 $0.095(0.07,0.12)$ $0.176(0.163,0.189)$
Unif(0,1) $+(1-2Bern(1/2))\cdot 0.1$ 1000 $0(-0.026,0.026)$ $0.176(0.163,0.189)$
Osteo model Perfect 1000 $0(-0.015,0.014)$ 0.054 (0.045,0.065)
Osteo model $+0.1$ 1000 $0.1(0.085,0.115)$ $0.064(0.057,0.073)$
Osteo model $+(1-2Bern(1/2))\cdot 0.1$ 1000 $0.03(0.015,0.045)$ $0.061(0.051,0.071)$
Table 3. Comparison of expectation of Brier score and quantiles under different simulation settings (Monte-Carlo simulation with $5000$ runs per setting)
Recall from (BIII) that the lowest expected Brier score occurs when predicted probabilities match the true probabilities, representing a perfect prediction that cannot be improved. However, we cannot observe the true probabilities. In some cases of observed incidences around $0.5$, perfect predictions yield a Brier score of close to 0.25, as shown in the following simulations. Hence if we observe a Brier score of around $\bar{y}-\bar{y}^{2}$, besides having a relatively bad model, the following alternative explanations could explain such a BS:
  • Many of the true probabilities are around $\bar{y}$ or $1-\bar{y}$, which can make the expected Brier score of perfect predictions close to $\bar{y}-\bar{y}^{2}$ (MII).
  • If $n$ is relatively low, as the Brier score is a realization of a random variable (BII), bigger random error are likely which can make the Brier score higher than its expectation.
Only if we can assume that neither of the above is excluded can conclude that the prediction model used is relatively bad.
As an illustration of example values, consider the median $\bar{y}-\bar{y}^{2}-BS_{perf}$ in Table 4, where $BS_{perf}$ is the Brier score under perfect predictions, which depends entirely on the underlying distribution, and again these medians strongly depend on the underyling distribution, but can be very low.
True Predictor n Median of $\bar{y}-\bar{y}^{2}-BS_{perf}$
distribution distribution (2,5%-&97,5%- Q)
Unif(0,1) Perfect 300 $0.083(0.059,0.105)$
Unif(0,1) Perfect 1000 $0.083(0.07,0.095)$
Beta(5,5) Perfect 300 $0.022(0.006,0.037)$
Beta(5,5) Perfect 1000 $0.022(0.014,0.031)$
Osteo model Perfect 300 $0.01(0.001,0.021)$
Osteo model Perfect 1000 $0.011(0.005,0.016)$
Smoke model Perfect 300 $0.026(0.01,0.041)$
Smoke model Perfect 1000 $0.026(0.017,0.034)$
Table 4.
While a priori we would not expect the Brier score for reasonable predictions to be bigger than $\bar{y}-\bar{y}^{2}$, it can happen. As we can not observe the true probabilities (BVI) we can not exclude the possibility that the expected Brier score for perfect predictions is close to the (unobservable) expected incidence $\bar{q}:=\frac{\sum_{i=1}^{n}q_{i}}{n}$, which is the expectation of $\bar{Y}$. Recall again from (BII) that the Brier score is a random variable. This randomness comes from the outcomes, making also the mean incidence $\bar{y}$ a random variable. All we know is that for perfect predictions, the probability to observe $BS>\bar{y}-\bar{y}^{2}$ will go to zero as $n\to\infty$ by (MIII). However, as in practice we will not be able to derive the perfect predictions, the expectation will be higher and this argument does not apply.
Hence as in the previous point, a Brier score above $\bar{y}-\bar{y}^{2}$ can be obtained when, apart from having a relatively bad model, the following alternative cases apply:
  • Many of true probabilities are around $0.5$, or tend to be close to each other.
  • If $n$ is relatively low, as the Brier score is a realization of a random variable (BII), random error can make the Brier score higher than its expectation by chance.
To illustrate this, we show that even for perfect predictions in some settings with low $n$, the probability of having a Brier score that is bigger than $\bar{y}-\bar{y}^{2}$ can be nonzero. As in practice we will not have perfect predictions, in the settings below for reasonable predictions the probability to have a Brier score bigger than $\bar{y}-\bar{y}^{2}$ is bigger, and the estimations can be understood as lower bounds for corresponding probabilities.
True Predictor n Event fraction
Distribution Distribution where $\bar{y}-\bar{y}^{2}>BS$
Unif(0,0.2) Perfect 300 0.060
Unif(0.3,0.7) Perfect 300 0.030
Osteroporosis model Perfect 300 0.019
Table 5. Comparison of expectation of Brier score and quantiles under different simulation settings (Monte-Carlo simulation with $5000$ runs per setting)
4. Connections to some other scores
We mention three other scores connected to BS. The Brier score from (1.1) equals mathematically the MSE, hence its square root is the root mean squared error (RMSE):
(4.1) $RMSE(p,y):=\sqrt{BS(p,y)}.$
(4.2) $MAE(p,y)=\frac{1}{n}\sum_{i=1}^{n}|p_{i}-y_{i}|,$
and the CIL [34]
(4.3) $CIL(p,y)=\frac{1}{n}\sum_{i=1}^{n}p_{i}-y_{i}.$
(4.4) $CIL(p,y)\leq MAE(p,y)\leq RMSE(p,y)$
as well as
(4.5) $MSE(p,y)\leq MAE(p,y)\leq RMSE(p,y).$
In particular, we note that both MAE and CIL are not strictly proper, which we also outline in the Appedix D.
5. Conclusions and final remarks
We adressed common misconceptions regarding the interpretation and use of the Brier score. Summarising, the observed Brier score is a function of
  1. (I)
    the underlying true probabilities,
  2. (II)
    the closeness of the predictions when compared toe the true probabilities,
  3. (III)
    some randomness that comes from the Bernoulli random variables.
In particular, the effect of the true underlying probabilities on the expectation of Brier score is strong, potentially stronger than the closeness of the predictions to the true probabilities as outlined in §$\S$ 2.1. We showed that a Brier score of zero is rather an indication of an error in realistic settings, and a very low Brier score does not necessarily indicate a perfect model. Indeed, the Brier score should be interpreted with caution, having in mind the relation to the true probabilities and the characteristics of the dataset. Direct comparisons of BSs across models on the same data can be done, and across different data should be avoided, or if done interpreted carefully. Additionally, we stressed that a low Brier score does not guarantee good calibration and should be complemented with additional calibration metrics. Furthermore the randomness in the Brier score, through the CLT, decreases as the sample size increases. A recent literature review on clinical prediction models [14] found that sample sizes used had median sample size $1250$ with $(Q1,Q3)=(353,188860)$. Hence at least a quarter of the prediction models had relatively low sample sizes, with randomness similar to the setting in our simulations with $n=300$.
Overall, while the Brier score remains a valuable metric for assessing probabilistic predictions, its interpretation requires an understanding of how the underlying probabilities and closeness of predictions influence the observed BS. Once misconceptions are avoided, the Brier score serves as a reliable relative measure of overall performance. In particular, it is strictly proper and hence in expectation, it is minimised uniquely at the true probabilities, making relative comparisons on the same data meaningful [9, Theorem 1], besides the nice order preservation property from (BIV).
Appendix A Why we care about the expectation: law of large numbers and central limit theorem

### v4

@@ A.1. Key properties about the Brier score
    Optimal predictions and true probabilities: The unique optimal prediction minimising the expected Brier score is the true outcome probability, i.e., for all $i$ $p_{i}=q_{i}$, as the Brier score is strictly proper [8, Theorem 1].
@@
    Expectation of Brier score:
    • $(n=1)$: As an illustrative example consider the case $n=1$. Using the true probability, we calculate the expectation. Let $Y_{1}\sim Bern(q_{1}),q_{1}\in[0,1]$ with prediction $p_{1}\in[0,1]$. Then the expectation of Brier score is
      (A.1) $g(p_{1},q_{1}):=\mathbb{E}_{Y_{1}\sim Bern(q_{1})}[BS(p_{1},Y_{1})]=p_{1}^{2}-2p_{1}q_{1}+q_{1}.$
      For the interested reader, the calculation is given in Appendix D.1. The optimal prediction minimising (A.1) is $p_{1}=q_{1}$. We next go through two key cases
@@
          if $q_{1}=1/2=p_{1}$, the expected Brier score is $f(1/2)=1/4$ corresponding to the maximum of (A.2),
@@
        The expectation of the Brier score as a function of $q_{1}$, when $p_{1}=q_{1}$, i.e., (A.2), looks as follows:
@@
          the expectation of the Brier score under perfect prediction but increased true probability, i.e., $\tilde{p}_{1}:=p_{1}+\varepsilon=\tilde{q}_{1}=q_{1}+\varepsilon$ for $q_{1}+\varepsilon\leq 1/2$, and taking the difference
          (A.3) $g(q_{1}+\varepsilon,q_{1}+\varepsilon)-g(q_{1},q_{1})=\varepsilon(1-2q_{1}-\varepsilon)$
          Note that the difference depends strongly on $q_{1}$, and that the same holds if we shift true probability and perfect prediction from $q_{1}>1/2$ to some value $q_{1}-\varepsilon$ for $q_{1}-\varepsilon\geq 1/2$.
@@
          (A.4) $g(q_{1}+\varepsilon,q_{1})-g(q_{1},q_{1})=\varepsilon^{2}$
          Note that the difference does not depend on $q_{1}$, and the same holds for $\tilde{p}_{1}:=p_{1}-\varepsilon$. The calculations are in Appendix §$\S$ D.2.
      • From the previous point, we conclude that if we slightly predict wrong, the costs are almost inexistent, e.g. if the true probability is $q_{1}$ but we predicted $p_{1}=q_{1}+0.1$, the expected difference between the Brier score of the perfect prediction and the slightly wrong one is $0.01$ by (A.4). However, if the true probability is changed towards $0.5$, e.g. from $q_{1}=0.1$ to $\tilde{q}_{1}=0.2$ and we have perfect prediction, the expected difference of the Brier scores is $0.07$ by (A.3) which is seven times more.
    • $n$ arbitrary: As the Brier score is a rescaled sum of the Brier score with $n=1$, the above conclusion extend in a straightforward way. We note the following nice order preservation property from (A.4): If $p_{1},\tilde{p_{1}}\in[0,1]$ such that $|p_{1}-q_{1}|<|\tilde{p_{1}}-q_{1}|$, then
@@
      This property is known as being effective [16], and for $n>1$ the order that is preserved is the order implied by the $l_{2}$ distance from $q\in[0,1]^{n}$ to the prediction $p\in[0,1]^{n}$.
@@
    Dependence of expectation and distribution of Brier score on true probabilities: The Brier score for multiple observations, as defined in (1.1), is the mean of the individual (one-dimensional) Brier scores. Its expectation and distribution depend on the true outcome probabilities $q_{i}$, which are generally unknown.
    As illustrated by (A.2) and Figure 5, the expected Brier score under perfect predictions varies with the distribution of the $q_{i}$. For example, if the $q_{i}$ are mostly concentrated near 0 or 1, the expected Brier score is close to 0. In contrast, if the $q_{i}$ cluster around 0.5, the expected Brier score under perfect predictions approaches 0.25.
@@
    Unobservability of true probabilities in clinical data: In practice, the true probability of an event occurring for an individual patient is not observable. Each patient is unique, and we only observe whether the event occurs or not (i.e., a binary outcome). This is in contrast to simulation settings where one can compare the true probabilities to estimated probabilities of regression or ML models, i.e. as done here [40, 12].
While individual-level probabilities will remain unknown, we can approximate their average value across a population by calculating the mean of observed outcomes overall, or in similar patient groups. This can be used to assess model calibration in clinical settings via calibration in the large (CIL) [6, 39]
(A.5) $CIL(p,y)=\frac{1}{n}\sum_{i=1}^{n}p_{i}-\frac{1}{n}\sum_{i=1}^{n}y_{i}.$
A.2. Notable mathematical features about the Brier score
Below, we summarise basic mathematical properties of the Brier score for the understanding of the reader. Mathematical details and arguments are given in Appendix D.3.
@@ A.2. Notable mathematical features about the Brier score
    Typical Brier score with perfect prediction for large $n$: Assume the true probabilities $q_{i}$ itself are realisations of random variables $Q_{i}\sim F$. Then, by the law of large numbers(LLN) the Brier score (1.1) for $n$ big roughly equals $\mathbb{E}[BS(Q_{1},Y_{1})]$, and probabilities for deviation from this value can be calculated via the central limit theorem (CLT). Hence $\mathbb{E}[BS(Q_{1},Y_{1})]$ roughly equals $BS((q_{1},\cdots,q_{n}),(y_{1},\cdots,y_{n}))]$ for $n$ large, and $\bar{y}$ roughly equals $\bar{q}$. Similarly if the predictions are given by $Q_{1}$ plus some iid error, the LLN and CLT appliy, and, e.g., tails can be analysed via large deviations theory. More detail on this perspective is in Appendix §$\S$ B.
Appendix B Why we care about the expectation: law of large numbers and central limit theorem